OpenStudy (jotopia34):

Calc 2 Help please... Does the series converge and if so to what? sum from n=1 to infinity of 1/ 1-(2/3)^n

OpenStudy (jotopia34):

\[\sum_{n=1}^{\infty}\frac{ 1 }{ 1-(\frac{ 2 }{ 3 } )^n}\]

OpenStudy (jotopia34):

I can't figure out how to take this limit. It looks like you would do 1/ 1-iinfinity. But how can I subtract infinity? Im lost